5. Complete the following table (next page) reflecting the total utility from either of two types of activities (per week): 1) “in-home” entertainment (friends visiting, etc.); and “on the town” entertainment (movies, restaurant meals, and the like). The price of in-home entertainment is $60. The price of a night on the town is $150.

In-Home Event

On-the-Town Event

#

times/week

Total Utility

Marginal Utility

MU/P

#

times/week

Total Utility

Marginal Utility

MU/P

0

0

0

0

1

400

1

800

2

780

2

1500

3

1100

3

1900

4

1380

4

2200

5

1500

5

2400

6

1600

6

2500

7

1650

7

2550

Now, imagine a “poor” graduate student with a weekly income of $330. What is her optimal allocation of spending between these two activities? What is her total utility with this allocation? What would be the graduate student’s allocation if the price of an “On-the-town” event fell to $60? What would be her total utility now? (Work with whole number quantities).
Answer #1

In-Home Event

On-the-Town Event

# Total Utility Marginal Utility

MU/P

(P=60)

# Total Utility Marginal Utility MU/P (P=150)

MU//P (P=60)

times/week times/week
0 0 0 0
1 400 400 6.66666667 1 800 800 5.33333333 13.33333
2 780 380 6.33333333 2 1500 700 4.66666667 11.66667
3 1100 320 5.33333333 3 1900 400 2.66666667 6.666667
4 1380 280 4.66666667 4 2200 300 2 5
5 1500 120 2 5 2400 200 1.33333333 3.333333
6 1600 100 1.66666667 6 2500 100 0.66666667 1.666667
7 1650 50 0.83333333 7 2550 50 0.33333333 0.833333

With the given income of $330,the student will purchase until her MU/P for In home = MU/P of on the town events

So,the student will purchase 3 units of In home Event and 1 unit of On the Town Event

Total utility = 1100+800 = 1900

When the price of On the town event falls to $60,then, the student will purchase

3 units of On the Town event and 2 units of In home Event

Total Utility = 780+1900 = 2680